5. At an activity level of 9,100 machine-hours in a month, Nooner Corporation's total variable production...

5.

At an activity level of 9,100 machine-hours in a month, Nooner Corporation's total variable production engineering cost is $803,530 and its total fixed production engineering cost is $161,280. What would be the total production engineering cost per unit, both fixed and variable, at an activity level of 9,600 machine-hours in a month? Assume that this level of activity is within the relevant range. (Do not round intermediate calculations.)
Answer #1

Variable cost per unit= $803530/9100= $88.30

Fixed cost per unit= $161280/9600= $16.80

Total production engineering cost per unit= Variable cost per unit+Fixed cost per unit

= $88.30+16.80= $105.10